# Data-Driven Global Sensitivity Analysis for Engineering Design Based on Individual Conditional Expectations

### v1

@@ 5.2.1 Global Sensitivity Analysis
Fig. 7(a) shows the GSA results of three metrics ($I_{pdp}$, $\mu_{I,ice}$, and $\bar{Sh}$) for the 5-variable wind turbine fatigue problem (the Sobol' indices are not shown due to the different scale, but instead shown in Fig. 8). All GSA metrics agree that the most important variable is $V_{hub}$ followed by $\theta_{w}$ and $H_{s}$. On the other hand, $T_{p}$ and $\theta_{p}$ are deemed to be insignificant. The averaged SHAP perceives the importance of $\theta_{w}$ is about half that of $H_{s}$, with $I_{pdp}$ and $\mu_{I,Ice}$ perceiving that the former is slightly important than the latter. It is interesting to see from the barplot that, although $V_{hub}$ is the most impactful variable, the effect of interaction is felt the strongest by $\theta_{w}$ and $H_{s}$ as indicated by the higher value of $\sigma_{I,ice}$ compared to $V_{hub}$. The cause for this will be discussed soon by visualizing the PDP and ICE curves. It is worth noting again that $\sigma_{I,ice}$ does not indicate which combinations of variables contribute the strongest to the interaction; however, the information from $\sigma_{I,ice}$ is useful in revealing the variability of the impact, complemented by the average impact from $\mu_{I,ice}$. Lastly, the impact of interactions on $T_{p}$ and $\theta_{p}$ is nevertheless small, and their importance can be neglected.
@@
The boxplots of ICE-based correlation values are shown in Fig. 7(b). The actual $\sigma_{\rho}$ values for $V_{hub}$, $\theta_{w}$, $H_{s}$, $T_{p}$, and $\theta_{p}$ are $0.058$, $0.275$, $0.160$, $0.444$, and $0.573$, respectively. These boxplots suggest that the interaction does not significantly alter the main trend of how $V_{hub}$ influences the moment. However, a notable shift in trends is evident for $\theta_{w}$ and $H_{s}$, indicating that the marginal behavior of these two variables is strongly influenced by their interactions with each other or with other variables. While the shifts in trend due to $T_{p}$ and $\theta_{p}$ appear significant, it is important to note that these changes may be the result of minor variations, as the impact of these two variables is relatively insignificant. Finally, quantification of the impact of interactions is shown in Fig. 8, with the diagonal entries showing the main effect (for Sobol' indices, the first-order Sobol' indices). All GSA agree that the most important interaction is between $H_{s}$ and $\theta_{w}$, followed by $V_{hub}-H_{s}$. Notice that the off-diagonal colors representing the correlations between $V_{hub}$, $\theta_{w}$, and $H_{s}$ are brighter for averaged SHAP and $I_{pdp}$ compared to Sobol' indices. This difference arises because Sobol' indices are variance-based, which amplifies the importance of the most influential variables.
(a) Barplot of $I_{pdp}$, $\mu_{I,ice}$, and $\bar{Sh}$
(b) Boxplot of ICE correlation values
@@
Figure 7: GSA results for the 5-variable wind turbine fatigue problem ($I_{pdp}$, $\mu_{I,ice}$, and $\bar{Sh}$) and the ICE correlation values.
(a) Sobol' indices
@@ 5.2.2 Visualization
Fig. 9 shows the PDPs and ICE curves for the three most significant variables, clearly showing how they affect the model’s behavior. The effect of $V_{hub}$ is nonlinear, with $M_{x,twr}$ generally increasing as $V_{hub}$ rises, although there is a slight decrease when $V_{hub}$ slightly exceeds its nominal value. Similarly, an increase in $\theta_{w}$ tends to raise $M_{x,twr}$. However, the actual trend influenced by $H_{s}$ is more complex, as revealed by the ICE plots and two-way PDPs. The harmonic looking pattern in $M_{x,twr}$ due to changes in $H_{s}$ is evident, with peaks observed around $H_{s}=-0.7$ and $H_{s}=0.7$. Additionally, as can be seen from the ICE plots, the impact on the bending moment is more pronounced when $\theta_{w}$ is large. Linking this observation with the GSA metrics, the large range of $\hat{f}_{pdp}$ of $V_{hub}$ contributes to its large importance. However, the impact of interaction itself is relatively small, as can be seen from the smaller deviation of ICE curves from the PDP. On the other hand, the deviation of ICE curves from their partial dependence function can be clearly seen for $\theta_{w}$ and $H_{s}$, particularly the latter; Overall, this contributes to their relatively large values of $\sigma_{I,ice}$. For comparison, Fig. 10 shows the SHAP dependence plots for the same set of variables. While the insights from SHAP and PDP/ICE are similar, the former appears more disordered, which is understandable given that SHAP values exist for every point in the input space. In that sense, it is easier to understand the effect of changing $V_{hub}$, $\theta_{w}$, and $H_{s}$ since PDP and ICE curves are smooth.
@@ 5.3.1 Data set
After pre-processing to remove erroneous aerodynamic data, particularly abnormal $C_{d}$$\alpha$ trends, the dataset was filtered, yielding 6,630 samples spanning 510 airfoils at 13 angles of attack. A snapshot of 100 airfoils with the airfoil created from the centre of the CST coefficients’ range (see Table 5) is depicted in Fig. 12(b). A data-driven PCE model with $p_{max}=4$ and $q=1$ was then constructed using 6,000 data points, with the remaining 630 points set aside for testing. The model achieved a coefficient of determination of 0.994, indicating a high level of accuracy suitable for knowledge discovery purposes.
@@ 5.3.2 Result
It can also be seen that the values of $\mu_{I,ice}$ are larger than $I_{pdp}$. In this regard, such differences are notably relatively significant for the first three lower surface CST variables, indicating that the respective $I_{pdp}$ values are smaller than $\mu_{I,ice}$ due to alternating positive and negative correlations (hence, an almost flat PDP, discussed next). It can also be seen that the impact of interactions between variables on the ICE curves is strong, as evidenced by the significantly high value of $\sigma_{I,ice}$ for all variables (see the error bar in Fig. 13). Therefore, the drag response surface is characterized by nonlinearity and strong interactions between input variables. It is worth noting that the values of $\sigma_{I,ice}$ themselves do not directly inform which variables contribute to the interactions for a specific input variable. Although it makes sense that the interaction is primarily due to $\alpha$ since it is the most important input variable, the interaction with other CST variables also contributes to such differences.
(a) GSA results
(b) ICE correlation values
@@
The boxplot of ICE correlation values presented in Fig. 13(b) offers several key insights. First, it is evident that lower surface CST parameters, particularly $A_{l,1}$ and $A_{l,4}$, exhibit high dispersion in terms of ICE correlation values, indicating significant variations in their ICE curves due to interactions with other variables (as confirmed in Fig. 15). Conversely, the lower dispersion for upper surface CST parameters suggests that interactions have a minimal effect on the overall trend, though the magnitude of the impact may still vary, as seen in the corresponding PDP and ICE plots. Overall, the lower surface CST parameters are more prone to trend changes due to interactions. In contrast, the upper surface parameters show less variation in trend but remain the most important variables in terms of significance. For the PDP and ICE curves of all variables, the trend consistently remains monotonic, reinforcing the hypothesis that the drag response of an airfoil is an unimodal function [53], even in the presence of nonlinearity.
The impact of $A_{l,1}$ is characterized by complex interaction with $A_{u,1}$ (see Fig. 15(a)). In this sense, a nonlinear trend that also changes in terms of correlation with the value of $A_{l,1}$ is present. Such an alternating trend results in a partial dependence function close to the flat line, which is why its $I_{pdp}$ is notably smaller than $\mu_{I,ice}$. The same trend also exists for $A_{l,4}$ as shown in Fig. 15(b). We decided to plot how $A_{l,4}$ interacts with $\alpha$ for demonstration purposes. In this regard, we can see that the association of the PDP of $A_{l,4}$ with its input value changes depending on the angle of attack. To be more exact, increasing $A_{l,4}$ results in higher drag for a positive angle of attack, while the opposite trend occurs for a smaller angle of attack. The PDP and ICE curves for $A_{u,2}$ clearly demonstrate its interaction with $A_{u,3}$ in influencing drag (see Fig. 15(c)). Notably, the effect of $A_{u,2}$ is most pronounced when $A_{u,3}$ is at a high value. Conversely, this implies that altering $A_{u,2}$ has minimal impact on drag when $A_{u,3}$ is small. Aerodynamically, the combinations of small $A_{u,2}$ and $A_{u,3}$ correspond to airfoils with reduced thickness, resulting in lower drag.
@@ 5.4 Interim summary
In the workflow of knowledge discovery in engineering, our ICE-based summaries $\mu_{I_{\mathrm{ice}},x_{j}}$ and $\sigma_{I_{\mathrm{ice}},x_{j}}$ are useful complements, since they capture instance-level spread and heterogeneity that scalar variance indices aggregate away. Based on the experiments that we performed, we summarize our main findings as follows:
@@
    The ICE-based feature importance ($\mu_{I,ice}$ with $\sigma_{I,ice}$) provides richer insights than the traditional PDP-based metric ($I_{pdp}$). While $\mu_{I,ice}$ captures effects missed by $I_{pdp}$, $\sigma_{I,ice}$ quantifies interaction influence. Complementing SHAP and Sobol’ indices, the ICE-based approach offers a distinct and informative perspective.
@@
    The value of $\sigma_{I,ice}$ reflects only the overall interaction impact, not which variables are involved. Interacting variables can be identified using joint PDPs, two-way Sobol’ indices, or two-way SHAP values. Combining these visualizations with GSA metrics ($\mu_{I,ice}$, $\sigma_{I,ice}$, and two-way GSA) enables a more comprehensive understanding of the model’s behavior.
@@
    The ICE correlation values provide a quantitative view of how interactions modify input–output relationships, revealing changes in linearity or correlation. However, they should be interpreted with caution, as they indicate association rather than the magnitude of impact.
@@
Our findings highlight the utility of ICE-based feature importance. While $\mu_{I,ice}$ offers a more comprehensive view of the input variable’s impact by capturing effects that $I_{pdp}$ might miss, $\sigma_{I,ice}$ further enriches the analysis by revealing how interactions with other variables influence this impact.
@@ 6 Conclusion and future works
The primary goal of this paper is to evaluate the effectiveness of PDP and ICE in supporting engineering design exploration and analysis. More importantly, we introduce new metrics that enhance the insights obtainable from existing PDP- and ICE-based approaches. Specifically, we propose an ICE-based feature importance metric that addresses the alternating-association issue inherent in PDP-based importance. The key idea is to compute the mean and standard deviation of the feature importance derived from each ICE curve, providing both a global average effect and information on how interactions with other variables influence a feature’s significance. This allows users to assess not only the average impact of an input variable but also its potential strongest or weakest influence on the predictive model. The standard deviation summarizes the effect of higher-order interactions in a single value. Additionally, we assess the correlation between ICE curves and PDP to further examine how interactions modify the input–output relationship.
@@
Future work may explore integrating PDP and ICE insights into optimization frameworks, using the proposed sensitivity metrics to guide design decisions by prioritizing influential variables. Advanced visualization techniques that dynamically reveal interactions captured by ICE curves and PDPs would further improve interpretability, especially in complex, nonlinear systems. The standard-deviation–based metrics could also support confidence-interval visualization for PDP, though a full uncertainty-quantification pipeline is needed to incorporate additional sources of uncertainty, including surrogate-model approximation errors. Developing such a pipeline would provide a more complete view of uncertainty in high-fidelity engineering simulations. Finally, automating insight extraction from PDP, ICE, and the proposed metrics could reduce reliance on expert judgment and improve usability in real-world engineering workflows.
Conflict of interest
The authors have no Conflict of interest to declare.

### v3

@@ 5.2.1 Global Sensitivity Analysis
Fig. 7(a) shows the GSA results of three metrics ($I_{pdp}$, $\mu_{I_{\mathrm{ice}}}$, and $\bar{Sh}$) for the 5-variable wind turbine fatigue problem (the Sobol’ indices are not shown due to the different scale, but instead shown in Fig. 8). All GSA metrics agree that the most important variable is $V_{hub}$ followed by $\theta_{w}$ and $H_{s}$. On the other hand, $T_{p}$ and $\theta_{p}$ are deemed to be insignificant. The averaged SHAP perceives the importance of $\theta_{w}$ is about half that of $H_{s}$, with $I_{pdp}$ and $\mu_{I_{\mathrm{ice}}}$ perceiving that the former is slightly more important than the latter. It is interesting to see from the barplot that, although $V_{hub}$ is the most impactful variable, the effect of interaction is felt the strongest by $\theta_{w}$ and $H_{s}$ as indicated by the higher value of $\sigma_{I_{\mathrm{ice}}}$ compared to $V_{hub}$. The cause for this will be discussed soon by visualizing the PDP and ICE curves. It is worth noting again that $\sigma_{I_{\mathrm{ice}}}$ does not indicate which combinations of variables contribute the strongest to the interaction; however, the information from $\sigma_{I_{\mathrm{ice}}}$ is useful in revealing the variability of the impact, complemented by the average impact from $\mu_{I_{\mathrm{ice}}}$. Lastly, the impact of interactions on $T_{p}$ and $\theta_{p}$ is nevertheless small, and their importance can be neglected.
@@
The violin plots of ICE-based correlation values are shown in Fig. 7(b). The actual $\sigma_{\rho}$ values for $V_{hub}$, $\theta_{w}$, $H_{s}$, $T_{p}$, and $\theta_{p}$ are $0.058$, $0.275$, $0.160$, $0.444$, and $0.573$, respectively. These violin plots suggest that the interaction does not significantly alter the main trend of how $V_{hub}$ influences the moment. However, a notable shift in trends is evident for $\theta_{w}$ and $H_{s}$, indicating that the marginal behavior of these two variables is strongly influenced by their interactions with each other or with other variables. While the shifts in trend due to $T_{p}$ and $\theta_{p}$ appear significant, it is important to note that these changes may be the result of minor variations, as the impact of these two variables is relatively insignificant. Finally, quantification of the impact of interactions is shown in Fig. 8, with the diagonal entries showing the main effect (for Sobol’ indices, the first-order Sobol’ indices). All GSA agree that the most important interaction is between $H_{s}$ and $\theta_{w}$, followed by $V_{hub}-H_{s}$. Notice that the off-diagonal colors representing the correlations between $V_{hub}$, $\theta_{w}$, and $H_{s}$ are brighter for averaged SHAP and $I_{pdp}$ compared to Sobol’ indices. This difference arises because Sobol’ indices are variance-based, which amplifies the importance of the most influential variables.
(a) Barplot of $I_{pdp}$, $\mu_{I_{\mathrm{ice}}}$, and $\bar{Sh}$
(b) Violin plot of ICE-based correlation values.
@@
Figure 7: GSA results for the 5-variable wind turbine fatigue problem ($I_{pdp}$, $\mu_{I_{\mathrm{ice}}}$, and $\bar{Sh}$) and the ICE correlation values.
(a) Sobol’ indices
@@ 5.2.2 Visualization
Fig. 9 shows the PDPs and ICE curves for the three most significant variables, clearly showing how they affect the model’s behavior. The effect of $V_{hub}$ is nonlinear, with $M_{x,twr}$ generally increasing as $V_{hub}$ rises, although there is a slight decrease when $V_{hub}$ slightly exceeds its nominal value. Similarly, an increase in $\theta_{w}$ tends to raise $M_{x,twr}$. However, the actual trend influenced by $H_{s}$ is more complex, as revealed by the ICE plots and two-way PDPs. The harmonic looking pattern in $M_{x,twr}$ due to changes in $H_{s}$ is evident, with peaks observed around $H_{s}=-0.7$ and $H_{s}=0.7$. Additionally, as can be seen from the ICE plots, the impact on the bending moment is more pronounced when $\theta_{w}$ is large. Linking this observation with the GSA metrics, the large range of $\hat{f}_{pdp}$ of $V_{hub}$ contributes to its large importance. However, the impact of interaction itself is relatively small, as can be seen from the smaller deviation of ICE curves from the PDP. On the other hand, the deviation of ICE curves from their partial dependence function can be clearly seen for $\theta_{w}$ and $H_{s}$, particularly the latter; Overall, this contributes to their relatively large values of $\sigma_{I_{\mathrm{ice}}}$. For comparison, Fig. 10 shows the SHAP dependence plots for the same set of variables. While the insights from SHAP and PDP/ICE are similar, the former appears more disordered, which is understandable given that SHAP values exist for every point in the input space. In that sense, it is easier to understand the effect of changing $V_{hub}$, $\theta_{w}$, and $H_{s}$ since PDP and ICE curves are smooth.
@@ 5.3.1 Data set
After removing erroneous aerodynamic data, particularly abnormal $C_{d}$$\alpha$ trends, the dataset comprised 6,600 samples from 510 airfoils at 13 angles of attack. Figure 12(b) shows 100 representative airfoils generated around the center of the CST coefficient range. A data-driven PCE model ($p_{\max}=4$, $q=1$) was trained on 6,000 samples and tested on 600, achieving $R^{2}=0.994$, demonstrating accuracy suitable for knowledge discovery.
@@ 5.3.2 Result
It can also be seen that the values of $\mu_{I_{\mathrm{ice}}}$ are larger than $I_{pdp}$. In this regard, such differences are notably relatively significant for the first three lower surface CST variables, indicating that the respective $I_{pdp}$ values are smaller than $\mu_{I_{\mathrm{ice}}}$ due to alternating positive and negative correlations. It can also be seen that the impact of interactions between variables on the ICE curves is strong, as evidenced by the significantly high value of $\sigma_{I_{\mathrm{ice}}}$ for all variables (see the error bar in Fig. 13). Therefore, the drag response surface is characterized by nonlinearity and strong interactions between input variables. It is worth noting that the values of $\sigma_{I_{\mathrm{ice}}}$ themselves do not directly inform which variables contribute to the interactions for a specific input variable. Although it makes sense that the interaction is primarily due to $\alpha$ since it is the most important input variable, the interaction with other CST variables also contributes to such differences.
(a) Barplot of $I_{pdp}$, $\mu_{I_{\mathrm{ice}}}$, and $\bar{Sh}$
(b) Violin plot of ICE-based correlation values.
@@
Variable $\sigma_{\rho}$
$\alpha$ 0.011
$A_{l,1}$ 0.705
$A_{l,2}$ 0.491
$A_{l,3}$ 0.309
$A_{l,4}$ 0.829
$A_{u,1}$ 0.319
$A_{u,2}$ 0.050
$A_{u,3}$ 0.012
$A_{u,4}$ 0.400
Table 6: Standard deviation of ICE correlation values for the 9-variable airfoil problem.
@@
The violin plot of ICE correlation values presented in Fig. 13(b) offers several key insights. The standard deviations of ICE correlation values are shown in Table 6. First, it is evident that lower surface CST parameters, particularly $A_{l,1}$ and $A_{l,4}$, exhibit high dispersion in terms of ICE correlation values, indicating significant variations in their ICE curves due to interactions with other variables (as confirmed in Fig. 15). Conversely, the lower dispersion for upper surface CST parameters suggests that interactions have a minimal effect on the overall trend, though the magnitude of the impact may still vary, as seen in the corresponding PDP and ICE plots. Overall, the lower surface CST parameters are more prone to trend changes due to interactions. In contrast, the upper surface parameters show less variation in trend but remain the most important variables in terms of significance. For the PDP and ICE curves of all variables, the trend consistently remains monotonic, reinforcing the hypothesis that the drag response of an airfoil is an unimodal function [6], even in the presence of nonlinearity.
The impact of $A_{l,1}$ is characterized by complex interaction with $A_{u,1}$ (see Fig. 15(a)). In this sense, a nonlinear trend that also changes in terms of correlation with the value of $A_{l,1}$ is present. Such an alternating trend results in a partial dependence function close to the flat line, which is why its $I_{pdp}$ is notably smaller than $\mu_{I_{\mathrm{ice}}}$. The same trend also exists for $A_{l,4}$ as shown in Fig. 15(b). We decided to plot how $A_{l,4}$ interacts with $\alpha$ for demonstration purposes. In this regard, we can see that the association of the PDP of $A_{l,4}$ with its input value changes depending on the angle of attack. To be more exact, increasing $A_{l,4}$ results in higher drag for a positive angle of attack, while the opposite trend occurs for a smaller angle of attack. The PDP and ICE curves for $A_{u,2}$ clearly demonstrate its interaction with $A_{u,3}$ in influencing drag (see Fig. 15(c)). Notably, the effect of $A_{u,2}$ is most pronounced when $A_{u,3}$ is at a high value. Conversely, this implies that altering $A_{u,2}$ has minimal impact on drag when $A_{u,3}$ is small. Aerodynamically, the combinations of small $A_{u,2}$ and $A_{u,3}$ correspond to airfoils with reduced thickness, resulting in lower drag.
@@ 5.4 Interim summary
In the workflow of knowledge discovery in engineering, our ICE-based summaries $\mu_{I_{\mathrm{ice}},x_{j}}$, $\sigma_{I_{\mathrm{ice}},x_{j}}$, and $\sigma_{\rho}$ are useful complements, since they capture instance-level spread and heterogeneity that scalar variance indices aggregate away. Based on the experiments that we performed, we summarize our main findings and key strengths of the proposed method as follows:
@@
    The ICE-based feature importance metrics ($\mu_{I_{\mathrm{ice}}}$ and $\sigma_{I_{\mathrm{ice}}}$) provide richer insights than the traditional PDP-based metric ($I_{pdp}$). While $\mu_{I_{\mathrm{ice}}}$ captures effects missed by $I_{pdp}$, $\sigma_{I_{\mathrm{ice}}}$ quantifies the heterogeneity of this influence across different conditional settings, thereby reflecting the strength of interaction effects. When used in conjunction with SHAP values and Sobol’ indices, the proposed ICE-based metrics offer a complementary and informative perspective that emphasizes conditional response behavior rather than variance attribution alone.
@@
    The value of $\sigma_{I_{\mathrm{ice}}}$ reflects only the overall interaction impact, not specifically on which variables are involved. Interacting variables can be identified using joint PDPs, two-way Sobol’ indices, or two-way SHAP values. Combining these visualizations with GSA metrics ($\mu_{I_{\mathrm{ice}}}$, $\sigma_{I_{\mathrm{ice}}}$, and two-way GSA) enables a more comprehensive understanding of the model’s behavior.
@@
    The ICE correlation values provide a quantitative view of how interactions modify input–output relationships, revealing changes in linearity or correlation. Their key strength lies in capturing directional and structural heterogeneity in conditional response behavior, which is not accessible through variance-based sensitivity measures alone. However, they should be interpreted with caution, as they indicate association rather than the magnitude of impact.
@@
Our findings highlight the utility of ICE-based feature importance. While $\mu_{I_{\mathrm{ice}}}$ offers a more comprehensive view of the input variable’s impact by capturing effects that $I_{pdp}$ might miss, $\sigma_{I_{\mathrm{ice}}}$ further enriches the analysis by revealing how interactions with other variables influence this impact. However, it is also important to discuss some limitations of the proposed metrics:
  1. 1.
    All proposed metrics (i.e., $\mu_{I_{\mathrm{ice}}}$, $\sigma_{I_{\mathrm{ice}}}$ and $\sigma_{\rho}$) are not suitable for handling correlated input variables. This limitation stems from the use of ICE itself, which evaluates conditional responses by varying one input at a time while sampling others independently. While this does not pose an issue in applications with independent inputs, such as design exploration or parametric studies, care must be taken when applying the proposed metrics to problems involving correlated variables.
  2. 2.
    The proposed $\sigma_{I_{\mathrm{ice}}}$, although useful to assess the impact of interactions of an input with other inputs, cannot distinguish which variables it strongly interacts with. Although visualization can help, it might be troublesome to look at every single other variable. It also lacks the capability to model higher-order interaction effects, such as third-order interactions. To that end, this analysis in terms of magnitude should be done together with the interaction Sobol’ indices.
  3. 3.
    The current method focuses only on global sensitivity in terms of how the input affects the output; it cannot be applied to specific cases, such as reliability-based sensitivity analysis, which focuses on extreme cases. This is in contrast to Sobol’ indices, which can be extended to reliability-based sensitivity analysis by using the indicator function in the calculation of the metrics [11].
The proposed metrics are model-agnostic and can be applied to any predictive model. When applied to a surrogate model constructed from data, however, the resulting sensitivity measures are evaluated with respect to the trained model and may therefore vary across different models approximating the same physical system. In such cases, the computed importance reflects the behavior learned by the surrogate model. Therefore, to accurately reflect the underlying modeled function, the surrogate itself must first achieve sufficient predictive accuracy.
@@ 6 Conclusion and future works
The primary goal of this paper is to evaluate the effectiveness of PDP and ICE in supporting engineering design exploration and analysis. Most importantly, we introduce new metrics that enhance the insights obtainable from existing PDP- and ICE-based approaches. Specifically, we propose ICE-based feature importance metrics that address the alternating-association issue inherent in PDP-based importance. The key idea is to compute the mean and standard deviation of the feature importance derived from each ICE curve, providing both a global average effect and information on how interactions with other variables influence a feature’s significance. This allows users to assess not only the average impact of an input variable but also its potential strongest or weakest influence on the predictive model. The standard deviation summarizes the effect of higher-order interactions in a single value. Additionally, we assess the correlation between ICE curves and PDP to further examine how interactions modify the input–output relationship.
@@
Despite these contributions, several limitations warrant discussion, and addressing them forms a central topic for future research. First, all proposed metrics ($\mu_{I_{\mathrm{ice}}}$, $\sigma_{I_{\mathrm{ice}}}$, and $\sigma_{\rho}$) inherit the limitations of ICE and are not suitable for correlated input variables. Second, while $\sigma_{I_{\mathrm{ice}}}$ summarizes the magnitude of interaction effects, it cannot identify which specific variables are responsible for strong interactions; thus, it should be complemented with interaction Sobol’ indices when detailed decomposition is needed. Third, the framework focuses on global sensitivity in terms of average input–output influence and is not directly applicable to reliability-based sensitivity analysis that emphasizes extreme events.
Future work may also explore integrating PDP and ICE insights into optimization frameworks, using the proposed sensitivity metrics to guide design decisions by prioritizing influential variables. Advanced visualization techniques that dynamically reveal interactions captured by ICE curves and PDPs would further improve interpretability, especially in complex, nonlinear systems. Incorporating uncertainty into the proposed metrics, especially uncertainty arising from surrogate-model approximation errors, necessitates a complete uncertainty-quantification pipeline. For probabilistic surrogate models, such as Gaussian process regression, predictive uncertainty information could be explicitly incorporated to quantify confidence in the extracted sensitivity measures. Developing such an integrated framework would provide a more comprehensive characterization of uncertainty in high-fidelity engineering simulations. Finally, automating insight extraction from PDPs, ICE curves, and the proposed metrics could reduce reliance on expert judgment and improve usability in real-world engineering workflows.
Acknowledgements
Pramudita Satria Palar would like to acknowledge financial support from Institut Teknologi Bandung through the Riset Kolaborasi Internasional 2024 scheme. This research was also funded by the Indonesian Endowment Fund for Education (LPDP) on behalf of the Indonesian Ministry of Higher Education, Science and Technology and managed under the EQUITY Program (Contract No. 8128/IT1.B07.1/TA.00/2025). This work is part of the activities of ONERA - ISAE - ENAC joint research group. The research presented in this paper has been performed in the framework of the COLOSSUS project (Collaborative System of Systems Exploration of Aviation Products, Services and Business Models) and has received funding from the European Union Horizon Europe program under grant agreement n 101097120 and in the MIMICO research project funded by the Agence Nationale de la Recherche (ANR) no ANR-24-CE23-0380.
Declaration of Competing interest
Pramudita Satria Palar reports that financial support was provided by Bandung Institute of Technology and LPDP. Nicolas Verstaevel and Paul Saves report that financial support was provided by the French National Research Agency. Joseph Morlier reports that financial support was provided by ISAE-SUPAERO Higher Institute of Aerospace Engineering. If there are other authors, they declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.